# Acting upon Imagination: When to Trust Imagined Trajectories in Model Based Reinforcement Learning

### v5

@@ 2 Related work
Lakshminarayanan et al. [14] utilized the ensembles of probabilistic NNs to determine predictive uncertainty. Kalweit and Boedecker [13] used a notion of uncertainty within a model-free RL (MFRL) agent to switch executing imagined trajectories from a dynamics model when MFRL agent has a high uncertainty. Conversely, Buckman et al. [2] used imagined trajectories to improve the sample complexity of a MFRL agent. They improved the Q function by using ensembles to estimate uncertainty and prioritize trajectories thereupon.
@@ 3 Preliminaries
Probabilistic Dynamics Model.
@@ 4 The Promise of Imagination
Trajectory Quality Analysis:
@@
Reward Analysis: Compared to the vector of state errors, the reward has the advantage of being a more compact representation (single scalar). It also provides substantial information. Figure 2 right shows the reward of successfully solved task in CP. After 50 steps, the reward does not change significantly and the system is at a local equilibrium. We contend that when the system is at equilibrium the dynamics model can reliably anticipate the outcomes of the agent’s actions, consequentially rewards are expected to remain similar $\Delta_{r}(t)=r_{t}-r_{t-1}\simeq\Delta_{r}(t+1)=r_{t+1}-r_{t}$.
@@ 5 Acting Upon Imagination
N-Skip
as a baseline, we introduce N-Skip, which is a straight forward method for replanning that executes a fixed $n$ steps in a trajectory (of length $H$) and triggers replanning at step $n+1$ ($n<H$). For $n=0$ the trajectory is recomputed at every step. As earlier replanning generally lead to lower error. $n$ is a hyperparameter that should be tuned to meet the performance requirements. In the CP environment, Figure 2 left shows a sharply increasing error at $n=7$, which amounts to $88\%$ less computations. Interestingly, despite its simplicity, N-Skip has not been extensively analyzed or reported in existing literature.
First Step Alike (FSA)
Some trajectories are more reliable than others and a cutoff of n skips for all trajectories does not consider this variation in quality. Figure 2 left shows that there are cases where after 19 steps of a trajectory, the prediction error is still lower than the average error of predicted states right after replanning. To account for such a variation, we propose a dynamic decision making.We would like to continue acting if the replanning will not improve over the error of predicted states. The error is lowest just after replanning and increases with number of steps. So, the main principle of FSA, is to omit replanning at any point $t+i,i<H$, as long as the error $\epsilon_{t+i-1}$ is comparable to errors right after replanning.
@@
Confidence Bounds (CB)
If the dynamics model has a notion of uncertainty, one can obtain a prediction $\hat{s}_{t+1}$ with an uncertainty or confidence $\sigma_{t+1}$. Given that $p^{*}_{s(t+1)}$ is modeled by an ensemble of Gaussian regressors, we can assume that the confidence bound $\sigma_{s(t+1)}$ represents the variability in predicted outcomes $\mu_{s(t+1)}$ of an action $a_{t}$, where the action in question has been deemed appropriate at the given state $s(t)$ therefore is in the trajectory. After performing the action, we obtain a real-world state $s_{t+1}$. This method considers the trajectory reliable, if the actual state $s_{t+1}$ is close to predicted state $\mu_{s(t+1)}$ within the confidence bound $\sigma_{s(t+1)}$ of predicted output states obtained from the dynamics model, meaning: $\mu_{s(t+1)}-c\times\sigma_{s(t+1)}\leq s_{t+1}\leq\mu_{s(t+1)}+c\times\sigma_{s(t+1)}$, where $c$ is a constant representing the selectivity of the filter adjusted with a factor of sigma, a hyper parameter to be tuned. The specific logic for the CB method is given by Algorithm 5. In a nutshell, this method assumes that the performance of an action could lead to several expected possible outcomes (bounded by the prediction). After performing the action, it is observed whether the outcome lies within the boundary of expected outcomes to determine the reliability of the trajectory.
@@
Probabilistic future trust (FUT)
FSA and CB asses the error of the state obtained after executing an action vs the prediction estimated by $\hat{f}_{\theta}$. Instead, FUT regards the effect of the last action on the outcomes of future actions, by projecting the remaining imagined actions in the trajectory from the newly obtained state. After replanning, the trajectory $\tau_{t+1}$ of H steps, where each step $i,i=0...H$ is composed of $\hat{s}_{t+i+1}$, $\sigma_{t+i+1}$, $\hat{r}_{t+i+1}$ sampled from $p^{*}_{s(t+i+1)}$ and $a^{*}_{t+i}$, offers a wealth of predicted information. FUT intends to detect whether after taking $a^{*}_{t+i}$ and reaching a new state $s_{t+i+1}\neq\hat{s}_{t+i+1}$ the rest of the predictions in $\tau_{t+i+1}$ still hold. Thus, we project the trajectory $\tau^{\prime}_{t+i+1}$ from state $s_{t+i+1}$ using imagined actions $a^{*}_{t+i:H}\in\tau_{t+i+1}$ and then we compare $\tau_{t+i+1}$ and $\tau^{\prime}_{t+i+1}$. If they differ, then the agent is deviated from the plan and it should trigger a replanning event. Otherwise, it proceeds to take action $a_{t+i+1}$. As long as the new trajectory is similar to the original estimation, we assume that the original plan is still valid and we skip that step. This does not mean that the optimal set of actions at each step is replanned. Rather, every time a replanning is skipped, we propagate only one trajectory starting from the current state of the simulator and still using the originally set of actions, $a^{*}$, as initially planned. Algorithm 6 describes FUT method. The original trajectory is estimated with the probability distribution $p^{*}_{s(t+1)}$ and the updated plan with $p^{\prime}_{s(t+1)}$. We use Kullback–Leibler divergence (KL) to evaluate the change in two distributions after each step in the simulator. We replan when there is larger difference than $\beta$ (a hyper-parameter). We control how far ahead the two distributions are compared by introducing a hyper-parameter: LA (look ahead steps).
@@
Bound Imagined Cost Horizon Omission (BICHO)
Figure 2 (right) shows that when the system reaches equilibrium, the reward stabilizes as well. BICHO assumes the expected reward is stable and attempts to determine whether deviations arise after each step in the trajectory.
@@ Appendix E EX-2: Online dynamics update
Figure 6, 7 and 8 show the resulting performance of different hyper parameters while training the dynamics model with skipping in CP, PU, RE and HC. Table 6 shows numerical results for each environment. CP was trained for 60 episodes, RE and PU for 150 and finally HC for 400. Rw represents the average over 3 experiments of the maximum rewards seen so far, RcPerMax is the percentage of replanning steps when the algorithm reached the maximum Rw. EpNr@@@@Max number of episodes needed to reach Rw and RelEp#@@@@Max represents the relative wall time compared to the baseline when the algorithm reached the maximum Rw.
Figure 6: Performance of different methods while training the dynamics model with skipping in the CP environment. Top) Episode reward (y-axis) in relation to the number of the relative wall time (x-axis). The vertical lines represent the end of training for the method in question. Middle) Percentage of replanning steps (y-axis) in relation to the number of training iterations (x-axis). Bottom) Mean error of the observations predicted by the dynamics model and the observation from the environment.
Figure 7: Performance of different methods while training the dynamics model with skipping in the PU environment. Top) Episode reward (y-axis) in relation to the number of the relative wall time (x-axis). The vertical lines represent the end of training for the method in question. Middle) Percentage of replanning steps (y-axis) in relation to the number of training iterations (x-axis). Bottom) Mean error of the observations predicted by the dynamics model and the observation from the environment.
Figure 8: Performance of different methods while training the dynamics model with skipping in the RE environment. Top) Episode reward (y-axis) in relation to the number of the relative wall time (x-axis). The vertical lines represent the end of training for the method in question. Middle) Percentage of replanning steps (y-axis) in relation to the number of training iterations (x-axis). Bottom) Mean error of the observations predicted by the dynamics model and the observation from the environment.
Figure 9: Performance of different methods while training the dynamics model with skipping in the HC environment. Top) Episode reward (y-axis) in relation to the number of the relative wall time (x-axis). The vertical lines represent the end of training for the method in question. Middle) Percentage of replanning steps (y-axis) in relation to the number of training iterations (x-axis). Bottom) Mean error of the observations predicted by the dynamics model and the observation from the environment.

### v6

@@ 2 Related work
Lakshminarayanan et al. [14] utilized the ensembles of probabilistic NNs to determine predictive uncertainty. Kalweit and Boedecker [13] used a notion of uncertainty within a model-free RL (MFRL) agent to switch executing imagined trajectories from a dynamics model when MFRL agent has a high uncertainty. Conversely, Buckman et al. [2] used imagined trajectories to improve the sample complexity of an MFRL agent. They improved the Q function by using ensembles to estimate uncertainty and prioritize trajectories thereupon.
@@ 3 Preliminaries
Probabilistic Dynamics Model
@@ 4 The Promise of Imagination
Trajectory Quality Analysis
@@
Reward Analysis. Compared to the vector of state errors, the reward has the advantage of being a more compact representation (single scalar). It also provides substantial information. Figure 2 right shows the reward of successfully solved task in CP. After 50 steps, the reward does not change significantly and the system is at a local equilibrium. We contend that when the system is at equilibrium the dynamics model can reliably anticipate the outcomes of the agent’s actions, consequentially rewards are expected to remain similar $\Delta_{r}(t)=r_{t}-r_{t-1}\simeq\Delta_{r}(t+1)=r_{t+1}-r_{t}$.
@@ 5 Acting Upon Imagination
N-Skip: as a baseline, we introduce N-Skip, which is a straight forward method for replanning that executes a fixed $n$ steps in a trajectory (of length $H$) and triggers replanning at step $n+1$ ($n<H$). For $n=0$ the trajectory is recomputed at every step. As earlier replanning generally lead to lower error. $n$ is a hyperparameter that should be tuned to meet the performance requirements. In the CP environment, Figure 2 left shows a sharply increasing error at $n=7$, which amounts to $88\%$ less computations. Interestingly, despite its simplicity, N-Skip has not been extensively analyzed or reported in existing literature.
First Step Alike (FSA): Some trajectories are more reliable than others and a cutoff of n skips for all trajectories does not consider this variation in quality. Figure 2 left shows that there are cases where after 19 steps of a trajectory, the prediction error is still lower than the average error of predicted states right after replanning. To account for such a variation, we propose a dynamic decision making. We would like to continue acting if the replanning will not improve over the error of predicted states. The error is lowest just after replanning and increases with number of steps. So, the main principle of FSA, is to omit replanning at any point $t+i,i<H$, as long as the error $\epsilon_{t+i-1}$ is comparable to errors right after replanning.
@@
Confidence Bounds (CB): If the dynamics model has a notion of uncertainty, one can obtain a prediction $\hat{s}_{t+1}$ with an uncertainty or confidence $\sigma_{t+1}$. Given that $p^{*}_{s(t+1)}$ is modeled by an ensemble of Gaussian regressors, we can assume that the confidence bound $\sigma_{s(t+1)}$ represents the variability in predicted outcomes $\mu_{s(t+1)}$ of an action $a_{t}$, where the action in question has been deemed appropriate at the given state $s(t)$ therefore is in the trajectory. After performing the action, we obtain a real-world state $s_{t+1}$. This method considers the trajectory reliable, if the actual state $s_{t+1}$ is close to predicted state $\mu_{s(t+1)}$ within the confidence bound $\sigma_{s(t+1)}$ of predicted output states obtained from the dynamics model, meaning: $\mu_{s(t+1)}-c\times\sigma_{s(t+1)}\leq s_{t+1}\leq\mu_{s(t+1)}+c\times\sigma_{s(t+1)}$, where $c$ is a constant representing the selectivity of the filter adjusted with a factor of sigma, a hyper parameter to be tuned. The specific logic for the CB method is given by Algorithm 5. In a nutshell, this method assumes that the performance of an action could lead to several expected possible outcomes (bounded by the prediction). After performing the action, it is observed whether the outcome lies within the boundary of expected outcomes to determine the reliability of the trajectory.
@@
Probabilistic future trust (FUT): FSA and CB asses the error of the state obtained after executing an action vs the prediction estimated by $\hat{f}_{\theta}$. Instead, FUT regards the effect of the last action on the outcomes of future actions, by projecting the remaining imagined actions in the trajectory from the newly obtained state. After replanning, the trajectory $\tau_{t+1}$ of H steps, where each step $i,i=0...H$ is composed of $\hat{s}_{t+i+1}$, $\sigma_{t+i+1}$, $\hat{r}_{t+i+1}$ sampled from $p^{*}_{s(t+i+1)}$ and $a^{*}_{t+i}$, offers a wealth of predicted information. FUT intends to detect whether after taking $a^{*}_{t+i}$ and reaching a new state $s_{t+i+1}\neq\hat{s}_{t+i+1}$ the rest of the predictions in $\tau_{t+i+1}$ still hold. Thus, we project the trajectory $\tau^{\prime}_{t+i+1}$ from state $s_{t+i+1}$ using imagined actions $a^{*}_{t+i:H}\in\tau_{t+i+1}$ and then we compare $\tau_{t+i+1}$ and $\tau^{\prime}_{t+i+1}$. If they differ, then the agent is deviated from the plan and it should trigger a replanning event. Otherwise, it proceeds to take action $a_{t+i+1}$. As long as the new trajectory is similar to the original estimation, we assume that the original plan is still valid and we skip that step. This does not mean that the optimal set of actions at each step is replanned. Rather, every time a replanning is skipped, we propagate only one trajectory starting from the current state of the simulator and still using the originally set of actions, $a^{*}$, as initially planned. Algorithm 6 describes FUT method. The original trajectory is estimated with the probability distribution $p^{*}_{s(t+1)}$ and the updated plan with $p^{\prime}_{s(t+1)}$. We use Kullback–Leibler divergence (KL) to evaluate the change in two distributions after each step in the simulator. We replan when there is larger difference than $\beta$ (a hyper-parameter). We control how far ahead the two distributions are compared by introducing a hyper-parameter: LA (look ahead steps).
@@
Bound Imagined Cost Horizon Omission (BICHO): Figure 2 (right) shows that when the system reaches equilibrium, the reward stabilizes as well. BICHO assumes the expected reward is stable and attempts to determine whether deviations arise after each step in the trajectory.
@@ Appendix E EX-2: Online dynamics update
Figure 6 and 7 show the resulting performance of different hyper parameters while training the dynamics model with skipping in CP, PU, RE and HC. Table 6 shows numerical results for each environment. CP was trained for 60 episodes, RE and PU for 150 and finally HC for 400. Rw represents the average over 3 experiments of the maximum rewards seen so far, RcPerMax is the percentage of replanning steps when the algorithm reached the maximum Rw. EpNr@@@@Max number of episodes needed to reach Rw and RelEp#@@@@Max represents the relative wall time compared to the baseline when the algorithm reached the maximum Rw.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 6: Performance of different methods while training the dynamics model with skipping. Top) Episode reward (y-axis) in relation to the number of the relative wall time (x-axis). The vertical lines represent the end of training for the method in question. Middle) Percentage of replanning steps (y-axis) in relation to the number of training iterations (x-axis). Bottom) Mean error of the observations predicted by the dynamics model and the observation from the environment. Left) Results in the CP environment. Right) Results in the PU environment.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 7: Performance of different methods while training the dynamics model with skipping. Top) Episode reward (y-axis) in relation to the number of the relative wall time (x-axis). The vertical lines represent the end of training for the method in question. Middle) Percentage of replanning steps (y-axis) in relation to the number of training iterations (x-axis). Bottom) Mean error of the observations predicted by the dynamics model and the observation from the environment. Left) Results in the RE environment. Right) Results in the HC environment.